# Colloid-like scale-free cluster-cluster aggregation during polymer collapse

### v1

@@ IV.2 Semi-flexible polymers
Figure 9: Deviation from the Viscek-Family-like scaling. Plot of $\Delta$ accounting the deviation from scaling relations in Eq.  (12) as a function of the bending stiffness for a polymer of length $N=2048$.
@@
Particularly for the scaling relation in Eq. (10), the exponent $\tau$ corresponding to the cluster size distribution becomes important. To extract the exponent, as in flexible polymer (see Fig. 5), we first verify if the dynamic scaling holds in semiflexible polymers as well by using the data for cluster-size distributions at different times during the collapse. Our exercise reveals that indeed the dynamic scaling holds for semiflexible polymers. In Figs. 8(a) and (b) we illustrate this finding for two $\kappa$ values for a polymer of length $N=2048$. There the plots of the scaling function $f(x)\equiv\langle s^{2}N_{s}(t)\rangle$ at different times as a function of the scaling variable $x\equiv s/t^{z}$ show reasonably good data collapse for both $\kappa$. For other values of $\kappa$ as well we obtain such data collapse. Note that the times used in (a) and (b) are different because as $\kappa$ increases the entry to the scaling regime gets delayed, as already discussed. In fact, for $\kappa=8$ one can notice that the data for $t=50$ and $100$ are a bit off for the same reason. In the region $s/t^{z}\ll 1$, the data show a power-law behavior with an expected exponent $\delta=2-\tau$, as discussed earlier. We observe consistent behavior in this regime, and extract $\delta$ and thereby $\tau$ from these plots by doing a fit using a power law
@@
in this regime. The results from the fitting exercise are quoted in Table 3. The results indicate that $\delta$ decreases monotonically as a function of $\kappa$, i.e., $\tau$ increases monotonically with $\kappa$. Now, recalling the estimated values of $z$ and $w$, reveals that as $\kappa$ increases one deviates from the relation in Eq. (12). The deviation can be quantified by $\Delta$ as a function of $\kappa$, where
@@
In cluster-cluster aggregation of purely particle systems, the validity of the scaling functions in Eqs. (8) or (11) depends on how the clusters diffuse before they coalescence with each other Meakin and Vicsek (1985). The diffusion constant of cluster of depends on the size $s$ as described in Eq. (13). As already pointed out, there exists a critical value of the exponent $\gamma_{c}$, such that for $\gamma>\gamma_{c}$ one realizes the scaling relation in Eq. (10) and for $\gamma<\gamma_{c}$ the relation in Eq. (12) holds true. In the present case, as in most of realistic situations, obviously $\gamma<0$, which indicates that $D_{s}$ should decrease with increasing $s$. In addition, our results for small $\kappa$ value also suggests that they belong to the scenario of $\gamma<\gamma_{c}$. To dig further about the deviation of the data from the scaling relation for $\kappa\geq 5$, one can rely on a comparative analyses of the diffusion constant $D_{s}$ as a function of $\kappa$.
@@
[…]
@@
The ratio of $R_{h}$ and $R_{g}$ gives a comparative idea about the shapes of the clusters of same size. The variation of $R_{h}$ with $\kappa$ is shown in in Fig. 10(a) for the two values of $s$. The profiles show a marginal dip around $\kappa=2$ following which it increases steadily. This suggests that clusters of same size will be more elongated and loosely bound as $\kappa$ increases, as also revealed from the typical clusters of size $s\approx 200$ for a small and large $\kappa$. Similarly, in Fig. 10(c) the ratio $R_{h}/R_{g}$ indicate a change in shape of the clusters around $\kappa\approx 5$. Smaller value of the ratio indicates a smaller fractal dimensions of the clusters with increasing $\kappa$. This the behavior of the data in Fig. 10 strongly implies a change in the diffusion constant $D_{s}$ with $\kappa$. A comparison with Fig. 9 confirms that the deviation from the scaling relation in Eq. (12) onsets around the same value of $\kappa$. Since the scaling relation in Eq. (12) is based on the diffusion of the clusters and its dependence on its size, this correlation is certainly strong. Here, we confirm that indeed there is a change in the diffusion of the clusters at around the same value of $\kappa$ where Eq. (12) breaks down. Of course, a detailed theoretical treatment of this correlation is needed to further substantiate it.
@@ V Discussion
The verification was performed by measuring the time dependence of the average cluster size $C_{s}(t)$, time dependence of the number of cluster of fixed size $N_{s}(t)$ and the cluster-size distribution at different times. Our results for the cluster growth indicates a universal power-law behavior $C_{s}(t)\sim t^{z}$ with $z\approx 1.65$, irrespective of $\kappa$. However, the exponent $w$ associated with the long-time decay $N_{s}(t)\sim t^{-w}$, varies inversely with $\kappa$. Interestingly, distributions of cluster sizes exhibit the dynamic scaling of the form $N_{s}(t)\sim s^{-2}f(s/t^{z})$, irrespective of $\kappa$. Otherwise difficult to estimate from our data, the realization of this scaling also allowed us to calculate the the exponent $\tau$ associated with the cluster-size distribution at a fixed time as $N_{s}\sim s^{-\tau}$. From the estimated values of $z$, $w$ and $\tau$ we have discovered that for small $\kappa<5$, in addition to the dynamic scaling, the relation $w=2z$ is also obeyed. For $\kappa\geq 5$ the estimated exponents show systematic deviation from the relation as a function of $\kappa$. This suggests that even though presence of dynamic scaling can be realized during polymer collapse, irrespective of the bending stiffness, the nature of scaling varies in terms of the relation between the dynamic exponents.
@@
The dynamic scaling, observed here for polymers is routinely observed for a wide range of colloidal aggregations. There, observation of a specific kind of scaling function is determined by factors such as size dependent diffusion constant of the clusters. In polymers, larger $\kappa$ lead to higher local order and different shape of the clusters compared to flexible or weakly stiff polymers. Thus, to find the origin of the breakdown of the scaling relation for large $\kappa$ in polymer collapse, we have estimated the hydrodynamic $R_{h}$ of the “pearls” or clusters. For the same monomer count $s$, the clusters belonging to semiflexible polymers show larger $R_{h}$, implying a smaller effective diffusion coefficient. The $\kappa$-dependent profile of $R_{h}$ follows the same trend as the scaling relation deviates from flexible polymers to semiflexible polymers. Thus, we speculate that the local order induced shape change as the cause of the change in the governing scaling law in cluster-cluster aggregation during polymer collapse.
@@
Historically, tracking the collapse of single polymers experimentally was difficult due to technical challenges in generating ultra-dilute solutions or obtaining sufficiently long monodisperse polymers Nishio et al. (1979); Chu et al. (1995). Significant progress in addressing these issues have been achieved through high resolution of modern instruments, e.g., small-angle X-ray scattering, single-molecule fluorescence, and dielectric spectroscopy have enabled probing of conformational changes in individual macromolecule with time Pollack et al. (2001); Sadqi et al. (2003); Tress et al. (2013); Hofmann et al. (2012); Naudi-Fabra et al. (2021). However, validation of the dynamic scaling behavior observed in this work would require well-resolved visualization of the intermediate states with clusters along the chain. While time-resolved investigations of colloidal aggregation and peptide assembly are now routinely performed Grober et al. (2023); Wu et al. (2017); Newcomb et al. (2012); Chakraborty et al. (2019), capturing the rapid collapse dynamics of polymers remains challenging due to spatio-temporal limitations of the conventional imaging techniques such as bright-field, confocal, time-resolved atomic force, or cryogenic transmission electron microscopy (cryo-TEM). Notably, the collapse time is expected to scale with polymer length as $\tau_{c}\sim N^{\nu}$, where typically $\nu=1-2$ Majumder et al. (2020). Thus, recent advances in synthesis of ultra-high-molecular-weight polymers have the potential to extend collapse times to the scale of hours Gao (2021); Chou et al. (2024); Eades et al. (2025). Additionally, viscosity and temperature of the ambient solvent can be tuned to prolong the collapse time Majumder et al. (2017, 2024). Under such conditions and in sufficiently dilute solutions, cryo-TEM may become a viable technique for imaging and quantifying the size distribution of intermediate clusters at various time points Newcomb et al. (2012); Patterson et al. (2017). Although such experimental efforts would require substantial resources and optimization, we anticipate that continued progress in polymer synthesis and imaging technologies will soon enable the direct observation of these transient collapse intermediates.
@@
For colloids, these scaling principles are known to be universal, independent of the composition of colloids Lin et al. (1989, 1990). The observation of similar principles in polymer expands the domain of the universality of these theories. The analogies with colloidal aggregations further reaffirms the validity of droplet coalescence model of polymer collapse. As a next step, it would be worth to verify the same for collapse of heteropolymers or protein-like heteropolymers. A step even further would be to perform all-atom MD simulations of polypeptide collapse and check the universality of the dynamic scaling. In this regard, one can also investigate the collapse of a real globular protein.

### v2

@@ IV.2 Semi-flexible polymers
Figure 9: Deviation from the Vicsek-Family-like scaling. Plot of $\Delta$ accounting the deviation from scaling relations in Eq. (12) as a function of the bending stiffness. For $N=8192$, results from only three different values of $\kappa$ are shown.
@@
Particularly for the scaling relation in Eq. (10), the exponent $\tau$ corresponding to the cluster size distribution becomes important. To extract the exponent, as in flexible polymer (see Fig. 5), we first verify if the dynamic scaling holds in semiflexible polymers as well by using the data for cluster-size distributions at different times during the collapse. Our exercise reveals that indeed the dynamic scaling holds for semiflexible polymers. In Figs. 8(a) and (b) we illustrate this finding for two $\kappa$ values for a polymer of length $N=2048$. There the plots of the scaling function $f(x)\equiv\langle s^{2}N_{s}(t)\rangle$ at different times as a function of the scaling variable $x\equiv s/t^{z}$ show reasonably good data collapse for both $\kappa$. For other values of $\kappa$ as well we obtain such data collapse. Note that the times used in (a) and (b) are different because as $\kappa$ increases the entry to the scaling regime gets delayed, as already discussed. In fact, for $\kappa=8$ one can notice that the data for $t=50$ and $100$ are a bit off for the same reason. In the region $s/t^{z}\ll 1$, the data show a power-law behavior with an expected exponent $\delta=2-\tau$, as discussed earlier. We observe a consistent behavior in this regime, and therefore consider the power law
@@
in this regime to calculate the average exponent as
$\langle\delta\rangle=\left\langle-\frac{d\ln f(x)}{d\ln x}\right\rangle,$ (23)
where the $\langle\dots\rangle$ now corresponds to a sliding $x$ window within the regime $x\ll 1$. The estimates obtained and the used $x$-range are quoted in Table 3. The results indicate that $\delta$ decreases monotonically as a function of $\kappa$, i.e., $\tau$ increases monotonically with $\kappa$. Now, recalling the estimated values of $z$ and $w$, reveals that as $\kappa$ increases one deviates from the relation in Eq. (12). The deviation can be quantified by $\Delta$ as a function of $\kappa$, where
@@
In cluster-cluster aggregation of purely particle systems, the validity of the scaling functions in Eqs. (8) or (11) depends on how the clusters diffuse before they coalescence with each other Meakin and Vicsek (1985). The diffusion constant of cluster depends on the size $s$ as described in Eq. (13). As already pointed out, there exists a critical value of the exponent $\gamma_{c}$, such that for $\gamma>\gamma_{c}$ one realizes the scaling relation in Eq. (10) and for $\gamma<\gamma_{c}$ the relation in Eq. (12) holds true. In the present case, as in most of realistic situations, obviously $\gamma<0$, which indicates that $D_{s}$ should decrease with increasing $s$. In addition, our results for small $\kappa$ value also suggests that they belong to the scenario of $\gamma<\gamma_{c}$. To dig further about the deviation of the data from the scaling relation for $\kappa\geq 5$, one can rely on a comparative analyses of the diffusion constant $D_{s}$ as a function of $\kappa$.
@@
where $d_{ij}=|\vec{r}_{i}-\vec{r}_{j}|$ is the distance between any two monomers $i$ and $j$ within the cluster of size $s$. Note that here $\langle\dots\rangle$ represents averaging over different clusters. Along the same line we also calculate the radius of gyration of the cluster as
@@
The ratio of $R_{h}$ and $R_{g}$ gives a comparative idea about the shapes of the clusters of same size. The variation of $R_{h}$ with $\kappa$ is shown in in Fig. 10(a) for the two values of $s$. The profiles show a marginal dip around $\kappa=2$ following which it increases steadily. This suggests that clusters of same size will be more elongated and loosely bound as $\kappa$ increases, as also revealed from the typical clusters of size $s\approx 200$ for a small and large $\kappa$. Similarly, in Fig. 10(b) the ratio $R_{h}/R_{g}$ indicate a change in shape of the clusters around $\kappa\approx 5$. Smaller value of the ratio indicates a smaller fractal dimensions of the clusters with increasing $\kappa$. This the behavior of the data in Fig. 10 strongly implies a change in the diffusion constant $D_{s}$ with $\kappa$. A comparison with Fig. 9 confirms that the deviation from the scaling relation in Eq. (12) onsets around the same value of $\kappa$. Since the scaling relation in Eq. (12) is based on the diffusion of the clusters and its dependence on its size, this correlation is certainly strong. Here, we confirm that indeed there is a change in the diffusion of the clusters at around the same value of $\kappa$ where Eq. (12) breaks down. Of course, a detailed theoretical treatment of this correlation is needed to further substantiate it.
Figure 11: Cluster growth for longer chins. (a) Cluster growth for flexible polymer with three different $N$. The inset shows the corresponding time dependence of the number of clusters $\langle n_{c}(t)\rangle$. (b) Corresponding variation of the cluster growth for three different $\kappa$ for the longest chain with $N=8192$. The inset presents the corresponding plots for $\langle n_{c}(t)\rangle$. The dashed lines in the main plots represent the same power law $\sim t^{1.67}$.
IV.3 Results for longer chains
Here we present results for a longer chain with $N=8192$ for a flexible polymer ($\kappa=0$) and two semiflexible polymers with $\kappa=5$ and $8$. The results presented here re-establish the claims made using relatively shorter chains. We start with presenting the cluster growth in Fig. 11. For a comparison, in Fig. 11(a) we show the data for the flexible polymer along with relatively shorter chains $N=4096$ and $2048$. Data for all system sizes follow each other until the shorter ones start deviating from the larger ones at late time before they reach the finite-size limit. This deviation is a virtue of the effect that at late times when the clusters are considerably huge, the number of clusters are limited and the tension in parts of the chain connecting them slowly becomes maximum. This eventually leads to a faster coalescence of the cluster than purely diffusive motion. This is in accordance with phenomenology described by Klushin Klushin (1998). Thus, for long enough chains the overall collapse possibly starts with the pearl-necklace phenomenology before crossing over to a chain-tension mediated coalescence, described by Klushin. A view of the full trajectory for longer chain in confirms this speculation. For small chain length, the duration of the latter stage is negligible before attaining the finite-size limit. The time dependence of the number of clusters, shown in the inset, also bears this signature with a steep fall at late times for longer chains.
Figure 12: Time dependence of the number of clusters of fixed sizes $N_{s}(t)$ for a long polymer with (a) $\kappa=0$, (b) $\kappa=5$ and (c) $\kappa=8$. The dashed lines show consistency of the data to the power-law decays obtained using a shorter chain.
The same observation can also be made from the cluster growths of polymers with $\kappa=5$ and $8$ for $N=8192$, as presented in Fig. 11(b). A careful observation of the trajectories for all three $\kappa$ will reveal that before the collapse reaches the late chain-tension mediated stage, the clusters are moving more or less in a Brownian path, resulting in a diffusive transport mechanism as observed in colloidal systems. However, the emergence of the late chain-tension mediated phase makes this diffusive scaling regime limited. For larger $\kappa$ one may expect that these regime may be further limited by the stiffness of the backbone and thereby effectively larger Kuhn lengths De Gennes (1979). However, since the considered $\kappa$ are in the moderate range ensuring that the final state is still a collapsed globule, this effect is perhaps not that magnified. On the other hand, within this moderate range the effective diffusive constant of the clusters decreases due to the increased $\kappa$, which affects only the amplitude of the cluster growth, and not the exponent as in other particle coarsening system Majumder and Das (2013). This smaller growth amplitude, consequently, prolongs the diffusive cluster-coalescence regime as evident from the plots in Fig. 11(b). Thus, results from polymer chains of length $N=2048$ were sufficient to focus on the cluster coalescence in the diffusive regime. In any case, for $N=8192$ also we have extracted the corresponding exponents within the scaling regime of interest.
For the cluster growth of the chain with $N=8192$, the obtained exponents are $\langle z\rangle=1.69(13)$, $1.67(19)$ and $1.64(02)$, respectively, for $\kappa=0$, $5$ and $8$, which is consistent with $z\approx 1.67$ quoted for the shorter chain with $N=2048$. Similarly, for the time-dependent decay of number of clusters of fixed size $N_{s}(t)$ are also consistent with the behavior claimed for a shorter chain, as shown in Fig. 12. The corresponding estimated values of $\langle w\rangle$ are $3.43(42)$, $2.69(06)$ and $2.33(02)$. Finally, in all the cases we have realized the dynamic scaling proposed in (8) and (11), as illustrated in Fig. 13. The corresponding estimated $\langle\delta\rangle$ are, respectively, $1.96(15)$, $1.60(17)$ and $1.57(07)$. Using these values of the exponents, we have also determined the deviation parameter $\Delta$ using Eq. (24). This calculated $\Delta$ for $N=8192$ for three values of $\kappa$ are also shown in Fig. 9, which within the error bars are in accordance with the data for $N=2048$.
Figure 13: Illustration of the dynamic scaling for a longer chain with (a) $\kappa=0$, (b) $\kappa=5$ and (c) $\kappa=8$. The dashed lines illustrate the consistency of the data with the respective power-laws obtained using a shorter chain in the regime $x\ll 1$.
@@ V Discussion
The verification was performed by measuring the time dependence of the average cluster size $C_{s}(t)$, time dependence of the number of cluster of fixed size $N_{s}(t)$ and the cluster-size distribution at different times. Our results for the cluster growth indicates a universal power-law behavior $C_{s}(t)\sim t^{z}$ with $z\approx 1.67$, irrespective of $\kappa$. However, the exponent $w$ associated with the long-time decay $N_{s}(t)\sim t^{-w}$, varies inversely with $\kappa$. Interestingly, distributions of cluster sizes exhibit the dynamic scaling of the form $N_{s}(t)\sim s^{-2}f(s/t^{z})$, irrespective of $\kappa$. Otherwise difficult to estimate from our data, the realization of this scaling also allowed us to calculate the the exponent $\tau$ associated with the cluster-size distribution at a fixed time as $N_{s}\sim s^{-\tau}$. From the estimated values of $z$, $w$ and $\tau$ we have discovered that for small $\kappa<5$, in addition to the dynamic scaling, the relation $w=2z$ is also obeyed. For $\kappa\geq 5$ the estimated exponents show systematic deviation from the relation as a function of $\kappa$. This suggests that even though presence of dynamic scaling can be realized during polymer collapse, irrespective of the bending stiffness, the nature of scaling varies in terms of the relation between the dynamic exponents.
@@
The dynamic scaling, observed here for polymers is routinely observed for a wide range of colloidal aggregations. There, observation of a specific kind of scaling function is determined by factors such as size dependent diffusion constant of the clusters. In polymers, larger $\kappa$ lead to higher local order and different shape of the clusters compared to flexible or weakly stiff polymers. Thus, to find the origin of the breakdown of the scaling relation for large $\kappa$ in polymer collapse, we have estimated the hydrodynamic $R_{h}$ of the “pearls” or clusters. For the same monomer count $s$, the clusters belonging to semiflexible polymers show larger $R_{h}$, implying a smaller effective diffusion coefficient. The $\kappa$-dependent profile of $R_{h}$ follows the same trend as the scaling relation deviates from flexible polymers to semiflexible polymers. Thus, we speculate that the local order induced shape change as the cause of the change in the governing scaling law in cluster-cluster aggregation during polymer collapse. To further confirm our findings we have also presented results for a polymer with a much longer chain length $N=8192$, albeit, for only three choices of $\kappa$. There, in addition to the pearl-necklace collapse, we found signature of a chain-tension mediated cluster growth at very late stage when the cluster are much larger. However, before that stage the cluster growth follows the same dynamic scaling as presented for the shorter chains. In future it would be interesting to investigate the scaling behavior of the final chain-tension mediated stage in detail, for which one needs to simulate even longer chains requiring more computational efforts. We propose this as a future task to explore.
@@
Historically, tracking the collapse of single polymers experimentally was difficult due to technical challenges in generating ultra-dilute solutions or obtaining sufficiently long mono-dispersed polymers Nishio et al. (1979); Chu et al. (1995). Significant progress in addressing these issues have been achieved through high resolution of modern instruments, e.g., small-angle X-ray scattering, single-molecule fluorescence, and dielectric spectroscopy have enabled probing of conformational changes in individual macromolecule with time Pollack et al. (2001); Sadqi et al. (2003); Tress et al. (2013); Hofmann et al. (2012); Naudi-Fabra et al. (2021). However, validation of the dynamic scaling behavior observed in this work would require well-resolved visualization of the intermediate states with clusters along the chain. While time-resolved investigations of colloidal aggregation and peptide assembly are now routinely performed Grober et al. (2023); Wu et al. (2017); Newcomb et al. (2012); Chakraborty et al. (2019), capturing the rapid collapse dynamics of polymers remains challenging due to spatio-temporal limitations of the conventional imaging techniques such as bright-field, confocal, time-resolved atomic force, or cryogenic transmission electron microscopy (cryo-TEM). Notably, the collapse time is expected to scale with polymer length as $\tau_{c}\sim N^{\nu}$, where typically $\nu=1-2$ Majumder et al. (2020). Thus, recent advances in synthesis of ultra-high-molecular-weight polymers have the potential to extend collapse times to the scale of hours Gao (2021); Chou et al. (2024); Eades et al. (2025). Additionally, viscosity and temperature of the ambient solvent can be tuned to prolong the collapse time Majumder et al. (2017, 2024). Under such conditions and in sufficiently dilute solutions, cryo-TEM may become a viable technique for imaging and quantifying the size distribution of intermediate clusters at various time points Newcomb et al. (2012); Patterson et al. (2017). Although such experimental efforts would require substantial resources and optimization, we anticipate that continued progress in polymer synthesis and imaging technologies will soon enable the direct observation of these transient collapse intermediates.
@@
For colloids, these scaling principles are known to be universal, independent of the composition of colloids Lin et al. (1989, 1990). The observation of similar principles in polymer expands the domain of the universality of these theories. The analogies with colloidal aggregations further reaffirms the validity of droplet coalescence model of polymer collapse. As a next step, it would be worth to verify the same for collapse of heteropolymers or protein-like heteropolymers. In this regard, for heteropolymers with a Gaussian and uniform distributions of polydispersity, the collapse dynamics is found to be universal, irrespective of the degree of polydispersity Singh and Shagolsem (2022). Albeit, it shows significant differences in the final monomer re-organization stage, which is crucial in achieving the final native state. A step even further would be to perform all-atom MD simulations of polypeptide collapse and check the universality of the dynamic scaling. In this regard, one can also investigate the collapse of a real globular protein.